Please write neat and clearly I can’t understand
General Chemistry Laboratory te: Lab 5 Naming Compounds and Writing Formulas Worksheet Student name Nhn PRE-LAOBRATORY QUESTIONS 1. Consider a mythical element X, which forms only two oxoanions: XO, and XO,2. Which of the two has a name ends in -ite? 2. Concerning the oxoanions in Question I, would the name of cither of them require a prefix such as hypo- or per-? Explain why or why not. 3. What is the role of Roman numerals in the names of the compounds formed by transitional metals? 4. Why do the names of the ionic compounds formed by group I and group 2 metals not include Roman numerals? IN-LABORATORY WORKSHEET
Answer #1

1) the - ite ending indicate low oxidation state

XO3 2-

x + 3*(-2) =-2

x = -2 + 6

x = +4

XO2 2-

x + 2*(-2 ) = -2

x = -2 + 4

x = 2

so XO2 2-   has name ends with -ite